OpenStudy (anonymous):

A number decreased by -3 is equal to -40. Write an equation.

OpenStudy (anonymous):

x-3=-40

OpenStudy (anonymous):

??

OpenStudy (anonymous):

A. -3 - n = - 40 B. n - 40 = -3 C. n - (-3) = -40 D. n + (-3) = - 40 those are the answers to choose from

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

decreased by -3 meaning -(-3)

OpenStudy (anonymous):

answer is c

OpenStudy (anonymous):

@NickR has it right. If it is confusing think about the difference of decreasing by three and decreasing by negative three. There has to be a difference between those statements.

OpenStudy (anonymous):

im going with a ?

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

glad to help